7. Hgb = 11g/dl Hct = 33% 8. Hgb = 6g/dl Het = 20% 9. Hgb = 15g/dl Hct = 45%

calculate the MCHC given the following information
Answer #1

Calculate MCHC by below given formula

MCHC=(Hgb g/dl x 100)÷HCT

Ans:

7) MCHC =(11×100)÷33

=1100÷33

=33.33 g/dl.

8) MCHC=(6×100)÷20

=600÷20

=30g/dl.

9)MCHC=(15×100)÷45

=1500÷45

=33.33g/dl.